OpenStudy (anonymous):

A cone with a radius of 8 feet and a slant height of 17 feet What is the surface area of this cone? Use 22/7 for pi

OpenStudy (anonymous):

SA = πr^2 + πrl

OpenStudy (anonymous):

can you explain it to me

OpenStudy (anonymous):

π = 22/7 r = radius l = slant height Just plug it in to the formula.

OpenStudy (anonymous):